# Tianxiang Chen USTC, China Alibaba Group, China

### 04 Mar 2024 v1

@@ 1 Introduction
In recent years, deep-learning-based methods have significantly improved ISTD performances, and most of them are CNN-based methods [10, 11, 12, 13, 14, 15, 16, 17]. However, the drawback of CNN-based methods is that their focus on local features is at the cost of global contexts. Global contexts are important to ISTD because the background pixels and small targets often seem so similar that cannot be distinguished by local features alone, which easily triggers missed detection. As a solution, some hybrid methods [18, 19, 20, 21, 22, 23] that combine ViT with CNN have been proposed to rely on ViT’s ability to model long-range dependencies. However, these methods generally suffer from a heavier computational burden due to the quadratic computational complexity of ViT. Despite certain work [19] adopting linear ViTs, its accuracy is still subordinate to the designs with quadratic complexity. Considering that high-resolution images are not rare in the infrared remote sensing domain (e.g. images produced by high-resolution infrared military sensors), this efficiency defect will be amplified, which hinders real-time ISTD when the resolution gets larger. How to relieve the inefficiency while maintaining high ISTD accuracy is our main concern.
@@
We aim to tailor a pure Mamba-based encoder to solve the locality defect while still maintaining a superior model efficiency. To this end, we get inspired by [29] and propose a novel Mamba-in-Mamba (MiM-ISTD) architecture for more efficient ISTD. To boost the feature representation ability of Mamba, we first divide the input images into several patches as “visual sentences” and then further divide them into sub-patches as “visual words”. Besides the outer Mamba blocks for extracting features of visual sentences, we further embed an inner Mamba block into the MiM block to excavate the local features of smaller visual words. In particular, features and relations between visual words in each visual sentence are calculated independently using a shared network to ensure the added amount of parameters and FLOPs is minimal. Then, the visual word features are consolidated back to their respective sentences. In this way, MiM-ISTD enables us to extract visual information with refined granularity. We present in Figure  1 that MiM-ISTD exhibits superior efficiency over other methods in terms of GPU memory usage and inference time, particularly as the resolution of infrared images increases. Also, MiM-ISTD can still exhibit notable accuracy-efficiency balance when processing infrared images of various resolutions.
@@
  • To the best of our knowledge, we are the first to apply Mamba to ISTD successfully, providing a new benchmark and some insights for the future advancements in efficient and potent Mamba-based methods.
@@ 3 Method
We illustrate the overall architecture of Mamba-in-Mamba (MiM) in Figure  3. It mainly consists of 1) a convolutional stem to produce patches and sub-patches, 2) a MiM pyramid encoder to extract both global and local features at different scales, 3) a decoder with several up-sample blocks to predict segmentation masks. We will describe our MiM in details in the following.
@@ 3.1 Preliminaries
State Space Models (SSMs) present significant challenges when applied to deep learning scenarios due to their continuous-time nature. To address this, it is imperative to transform the ODE into a discrete function through a discretization process. The transformation is essential for synchronizing the model with the sample rate of the underlying signal embodied in the input data, facilitating computationally efficient operations [16]. Considering the input $x_{k}\in\mathbb{R}^{L\times D}$, a sampled vector within the signal flow of length $L$ following [17], a timescale parameter $\Delta$ can be introduced to the continuous parameters $A$ and $B$ into their discrete counterparts $\overline{A}$ and $\overline{B}$ following the zeroth-order hold (ZOH) rule. Consequently, Equation  1 can be discretized as follows:
@@
where $B,C\in\mathbb{R}^{D\times N}$ and $\Delta\in\mathbb{R}^{D}$. In practice, following [12], we refine the approximation of $\overline{B}$ using the first-order Taylor series:
@@
The distinction between 2D visual data and 1D language sequences makes it unsuitable to directly apply Mamba to vision tasks. For instance, while 2D spatial information is vital in vision tasks, it is not the primary focus in 1D sequence modeling, resulting in restricted receptive fields that overlook potential relationships with patches unscanned. The 2D selective scan (SS2D) in [27] can address the issue. The overview of SS2D is depicted in Figure  2. SS2D arranges image patches in four different directions to generate four separate sequences. The quad-directional scanning strategy ensures that each element in the feature map integrates information from all other locations in different directions, thus creating a global receptive field without increasing linear computational complexity. Each resulting feature sequence is then processed using the selective scan space state sequential model (S6). The feature sequences are combined to reconstruct the 2D feature map. Given the input feature $z$, then the output feature $\overline{z}$ of SS2D can be expressed as:
@@
Figure 3: Overview of our MiM-ISTD, which mainly includes a convolutional stem, a pure Mamba-based MiM pyramid encoder and a plain decoder.
@@ 3.2 Mamba-in-Mamba for Efficient ISTD
Present ISTD methods mainly typically employ CNNs or hybrids of CNNs and ViTs. The latter compensate for the shortcomings of the former in modeling long-range dependencies, but suffer from a quadratic computational complexity. Recently, Mamba has been proposed. It is renowned for its superior model efficiency, less GPU memory usage and better long-range dependency modelling and has been applied to the vision domain. Therefore, we try to explore whether Mamba can also be applied to improve ISTD performances. However, when we directly apply vision mamba block to ISTD, the accuracy is not very impressive because local features, which matter a lot to detect small targets, are less explored. To address this deficiency, we propose a Mamba-in-Mamba (MiM-ISTD) architecture, shown in Figure  3, to learn both global and local information in an image while guaranteeing superb model efficiency.
Given a 2D infrared image $\mathcal{X}\in\mathbb{R}^{H\times W\times 1}$, it is divided evenly into $n$ patches to form $\mathcal{X}=[X^{1},X^{2},...,X^{n}]$ where each patch is in $\mathbb{R}^{n\times p\times p\times 1}$, with $(p,p)$ denoting the resolution of each patch. In MiM-ISTD, we view the patches as visual sentences that represent the image. Further, each patch is segmented into $m$ smaller sub-patches, making each visual sentence a sequence of visual words:
@@
where $x^{i,j}\in\mathbb{R}^{s\times s\times 1}$ is the $j-th$ visual word of the $i-th$ visual sentence, $(s,s)$ is the spatial size of sub-patches. Since our MiM adopts a pyramid encoder structure, the spatial shapes of visual sentences and words are unfixed and will gradually decrease as the network layers deepen.
@@
We construct a convolutional stem, where a stack of $3\times 3$ convolutions is utilized, to produce visual words $\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times C}$ and visual sentences $\in\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times D}$ at the first stage, where $C$ is the visual word dimension and $D$ is the visual sentence dimension. Each visual word corresponds to a $4\times 4$ pixel region in the original image, and each visual sentence is composed of $4\times 4$ visual words. Unlike ViTs, we do not use add the position embedding bias to visual words and sentences due to the causal nature of visual mamba block [27].
3.2.2 MiM Pyramid Encoder.
The core part of our MiM is its pyramid encoder of four stages with different number of tokens, as shown in Figure  3. Across the four stages, the spatial shape of visual words are set as $\frac{H}{2}\times\frac{W}{2}$, $\frac{H}{4}\times\frac{W}{4}$, $\frac{H}{8}\times\frac{W}{8}$ and $\frac{H}{16}\times\frac{W}{16}$. The spatial shape of visual sentences are set as $\frac{H}{8}\times\frac{W}{8}$, $\frac{H}{16}\times\frac{W}{16}$, $\frac{H}{32}\times\frac{W}{32}$ and $\frac{H}{64}\times\frac{W}{64}$. We adopt the patch merging in [38] as the down-sampling operation. Each stage consists of multiple MiM blocks which process both word-level and sentence-level features. The Visual words are mapped to a sequence of word embeddings via a linear projection, expressed as:
@@
where $w^{i,j}\in\mathbb{R}^{c}$ is the $j-th$ word embedding of the $i-th$ visual sentence, $c$ is the dimension of word embedding, and $Vec(·)$ refers to the vectorization operation.
@@
where $l$ is the index of the $l-th$ block, and $L$ is the total count of such blocks. The first word embedding input $W^{i}_{0}$ is the $W^{i}$ in Equation  6. All transformed word embeddings are denoted by $\mathcal{W}_{l}=[W^{1}_{l},W^{2}_{l},...,W^{n}_{l}]$. This can be viewed as an inner Mamba block, illustrated in Figure  4 (a). In this process, the relationships among visual words is built by computing interactions among each two visual words. For example, in a patch containing a small target, a word denoting the target would have a stronger relation with other target-related words while interact less with the background part.
At the sentence level, we generate sentence embedding memories as storage for the sequence of sentence-level representations, initialized as zero: $\mathcal{S}_{0}=[S^{1}_{0},S^{2}_{0},...,S^{n}_{0}]\in\mathbb{R}^{n\times d}$. In each layer, the sequence of word embeddings are mapped to the domain of sentence embedding by linear projection and subsequently integrated into the sentence embedding:
@@
The outer Mamba block can model the relationships among sentence embeddings. The inputs and outputs of the MiM block are the visual word embeddings and sentence embeddings, so the MiM pyramid encoder can be defined as
@@
In our MiM block, both the inner Mamba block and the outer Mamba block adopt the visual mamba block [27], as shown in Figure  4 (b). The inner Mamba block models the relationship between visual words for local feature extraction, while the outer Mamba block captures the global feature by modelling the relationship between visual sentences. We set $L_{1},L_{2},L_{3},L_{4}=2,4,9,2$ by default. Stacking the MiM blocks for $L=17$ layers, we build our MiM pyramid encoder.
We feed the visual sentence embedding outputs of each stages to respective decoder stages. The spatial shapes of the four stage encoder outputs are $\frac{H}{8}\times\frac{W}{8}$, $\frac{H}{16}\times\frac{W}{16}$, $\frac{H}{32}\times\frac{W}{32}$ and $\frac{H}{64}\times\frac{W}{64}$, which are different from the feature map scales from typical backbones. To solve this discrepancy, these outputs should pass through straightforward up-sampling modules before reaching the decoder. The module includes a $2\times 2$ transposed convolution with a stride of two, succeeded by batch normalization [44], GeLU [45], a stride-one $3\times 3$ convolution, another batch normalization and GeLU. Therefore our MiM pyramid encoder can produce feature maps with strides of 4, 8, 16, and 32 pixels relative to the input image.
@@ 4.1 Experimental Settings
where $A_{i}$ and $A_{u}$ are the areas of intersection region and union region between the prediction and ground truth, respectively. $N$ is the total number of samples, $TP[.]$ is the number of true positive pixels, $T[.]$ and $P[.]$ is the number of ground truth and predicted positive pixels.
Specifically, $nIoU$ is the normalized $IoU$, which can make a better balance between structural similarity and pixel accuracy of infrared small targets. $P_{d}$ calculates the proportion between the number of correctly predicted targets $N_{pred}$ and all targets $N_{all}$. $F_{a}$ refers to the ratio of falsely predicted target pixels $N_{false}$ and all the pixels in the infrared image $N_{all}$. $P_{d}$ and $F_{a}$ are calculated as follows:
@@ 4.3 Visualization
4.4 Ablation study
4.4.1 Ablation study of Each Module.
4.4.2 Ablation Study of the Granularity of Patch Division.
To explore the impact of information contained in each visual word, we adjust the representation granularity of sub-patches so that each visual word corresponds to a $2\times 2$ pixel region in the original image (No.#$\#1$), in contrast to $4\times 4$ of our present setting. We also change the initial spatial shape of visual sentence from $\frac{H}{8}\times\frac{W}{8}$ to $\frac{H}{4}\times\frac{W}{4}$ (No.#$\#2$) to ablate the effect of visual sentence granularity. We adopt the configuration that demonstrates the best performance as our present setting.
@@ 4.4 Ablation study
##\##2 Visual Sentence $\rightarrow H/4\times W/4$ 79.67 78.04 100 3.685
##\##3 Ours 80.80 80.20 100 2.179

### 08 Mar 2024 v2

@@ 1 Introduction
In recent years, deep-learning-based methods have significantly improved ISTD performances, and most of them are CNN-based methods [12, 13, 14, 15, 16, 17, 10, 18]. However, the drawback of CNN-based methods is that their focus on local features is at the cost of global contexts. Global contexts are important to ISTD because the background pixels and small targets often seem so similar that cannot be distinguished by local features alone, which easily triggers missed detection. As a solution, some hybrid methods [11, 19, 20, 21, 22, 23, 24] that combine ViT with CNN have been proposed to rely on ViT’s ability to model long-range dependencies. However, these methods generally suffer from a heavier computational burden due to the quadratic computational complexity of ViT. Despite certain work [19] adopting linear ViTs, its accuracy is still subordinate to the designs with quadratic complexity. Considering that high-resolution images are not rare in the infrared remote sensing domain (e.g. images produced by high-resolution infrared military sensors), this efficiency defect will be amplified when the resolution gets larger and hinder real-time ISTD. How to relieve the inefficiency while maintaining high ISTD accuracy is our main concern.
@@
We aim to tailor a pure Mamba-based encoder to solve the locality defect while still maintaining a superior model efficiency. To this end, we get inspired by TNT [30], which effectively models local structural features with a trivial increase of computation and memory cost, and propose a novel Mamba-in-Mamba (MiM-ISTD) architecture for more efficient ISTD. To boost the feature representation ability of Mamba, we divide the input image into several patches as "visual sentences" and then further divide them into sub-patches as "visual words". We use an outer Mamba block to extract features of visual sentences, and further assist it with inner Mamba blocks to excavate the local features of smaller visual words. In particular, features and relations between visual words in each visual sentence are calculated independently using a shared network to ensure the added amount of parameters and FLOPs is minimal. Then, the visual word features are consolidated back into their respective sentences. In this way, MiM-ISTD enables us to extract visual information with refined granularity. We present in Figure  1 that MiM-ISTD exhibits superior efficiency over other methods in terms of GPU memory usage and inference time, particularly as the resolution of infrared images increases. In general, MiM-ISTD can achieve the most notable accuracy-efficiency balance compared with other SOTA methods.
@@
  • To the best of our knowledge, we are the first to apply Mamba to ISTD successfully, providing a new benchmark and valuable insights for future advancements in efficient and potent Mamba-based methods.
@@ 3 Method
In this section, we describe our proposed Mamba-in-Mamba for efficient ISTD and analyze the computation complexity in detail.
@@ 3.1 Preliminaries
State Space Models (SSMs) present significant challenges when applied to deep learning scenarios due to their continuous-time nature. To address this, the ODE needs to be transformed into a discrete function. Considering the input $x_{k}\in\mathbb{R}^{L\times D}$, a sampled vector within the signal flow of length $L$ following [45], a timescale parameter $\Delta$ can be introduced to the continuous parameters $A$ and $B$ into their discrete counterparts $\overline{A}$ and $\overline{B}$ following the zeroth-order hold (ZOH) rule. Consequently, Equation  1 can be discretized as follows:
@@
where $B,C\in\mathbb{R}^{D\times N}$ and $\Delta\in\mathbb{R}^{D}$. In practice, we refine the approximation of $\overline{B}$ using the first-order Taylor series:
@@
2D spatial information cannot be effectively captured by models designed for 1D data, making it unsuitable to directly apply Mamba to vision tasks. The 2D selective scan (SS2D) in [28] can address the issue. The overview of SS2D is depicted in Figure  2. SS2D arranges image patches in four different directions to generate four separate sequences. The quad-directional scanning strategy ensures that each element in the feature map integrates information from all other locations in different directions, thus creating a global receptive field without increasing linear computational complexity. Each resulting feature sequence is then processed using the selective scan space state sequential model (S6) before merging to reconstruct the 2D feature map. Given the input feature $z$, then the output feature $\overline{z}$ of SS2D can be expressed as:
@@
Figure 3: Overview of our MiM-ISTD, which mainly includes a convolutional stem, a pure Mamba-based MiM hierarchical encoder, and a plain decoder.
@@ 3.2 Mamba-in-Mamba for Efficient ISTD
Present ISTD methods mainly typically employ CNNs or hybrids of CNNs and ViTs. The latter compensates for the shortcomings of the former in modeling long-range dependencies but suffers from a quadratic computational complexity. Recently, Mamba has been proposed. It is renowned for its superior model efficiency, less GPU memory usage, and better long-range dependency modeling and has been successfully applied to the vision domain. Therefore, we explore whether Mamba can also be applied to improve ISTD performances. However, when we directly apply visual mamba block to ISTD, the accuracy is not very impressive because local features, which matter a lot to detect small targets, are less explored. To address this deficiency, we propose a Mamba-in-Mamba (MiM-ISTD) architecture, shown in Figure  3, to learn both global and local information in an image while guaranteeing superb model efficiency.
Given a 2D infrared image $\mathcal{X}\in\mathbb{R}^{H\times W\times 3}$, it is divided evenly into $n$ patches to form $\mathcal{X}=[X^{1},X^{2},...,X^{n}]$ where each patch is in $\mathbb{R}^{n\times p\times p\times 3}$, with $(p,p)$ denoting the resolution of each patch. In MiM-ISTD, we view the patches as visual sentences that represent the image. Further, each patch is segmented into $m$ smaller sub-patches, making each visual sentence a sequence of visual words:
@@
where $x^{i,j}\in\mathbb{R}^{s\times s\times 3}$ is the $j-th$ visual word of the $i-th$ visual sentence $X^{i}$, $(s,s)$ is the spatial size of sub-patches. Since our MiM adopts a hierarchical encoder structure, the spatial shapes of visual sentences and words are unfixed and will gradually decrease as the network layers deepen.
@@
We construct a convolutional stem, where a stack of $3\times 3$ convolutions is utilized, to produce visual words $\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times C}$ and visual sentences $\in\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times D}$ at the first stage, where $C$ is the visual word dimension and $D$ is the visual sentence dimension. Each visual word corresponds to a $4\times 4$ pixel region in the original image, and each visual sentence is composed of $4\times 4$ visual words. Unlike ViTs, we do not add the position embedding bias to visual words and sentences due to the causal nature of visual mamba block [28].
3.2.2 MiM Hierarchical Encoder.
The core part of our MiM is its hierarchical encoder of four stages with different numbers of tokens, as shown in Figure  3. Across the four stages, the spatial shape of visual words is set as $\frac{H}{2}\times\frac{W}{2}$, $\frac{H}{4}\times\frac{W}{4}$, $\frac{H}{8}\times\frac{W}{8}$ and $\frac{H}{16}\times\frac{W}{16}$. The spatial shape of visual sentences are set as $\frac{H}{8}\times\frac{W}{8}$, $\frac{H}{16}\times\frac{W}{16}$, $\frac{H}{32}\times\frac{W}{32}$ and $\frac{H}{64}\times\frac{W}{64}$. We adopt the patch merging in [39] as the down-sampling operation. Each stage consists of multiple MiM blocks which process both word-level and sentence-level features. The visual words $x^{i,j}$ are mapped to a sequence of word embeddings $w^{i,j}$ via a linear projection, expressed as:
@@
where $w^{i,j}\in\mathbb{R}^{c}$ is the $j-th$ word embedding of the $i-th$ visual sentence, $c$ is the dimension of word embedding, $W^{i}$ is the collection of word embeddings of the $i-th$ visual sentence, and $Vec(·)$ refers to the vectorization operation.
@@
where $l$ is the index of the $l-th$ block, and $L$ is the total count of such blocks. The first word embedding input $W^{i}_{0}$ is the $W^{i}$ in Equation  6. All transformed word embeddings are denoted by $\mathcal{W}_{l}=[W^{1}_{l},W^{2}_{l},...,W^{n}_{l}]$. This can be viewed as an inner Mamba block, illustrated in Figure  4 (a). In this process, the relationships among visual words are built by computing interactions among each two visual words. For example, in a patch containing a small target, a word denoting the target would have a stronger relation with other target-related words while interacting less with the background part.
At the sentence level, we generate sentence embedding memories as storage for the sequence of sentence-level representations, initialized as zero: $\mathcal{S}_{0}=[S^{1}_{0},S^{2}_{0},...,S^{n}_{0}]\in\mathbb{R}^{n\times d}$. In each layer, the sequence of word embeddings is mapped to the domain of sentence embedding by linear projection and subsequently integrated into the sentence embedding:
@@
The outer Mamba block can model the relationships among sentence embeddings. The inputs and outputs of the MiM block are visual word embeddings and sentence embeddings, so the MiM hierarchical encoder can be defined as
@@
In our MiM block, both the inner Mamba block and the outer Mamba block adopt the visual mamba block [28], as shown in Figure  4 (b). The inner Mamba block models the relationship between visual words for local feature extraction, while the outer Mamba block captures the global feature by modelling the relationship between visual sentences. We set the block number $L_{1},L_{2},L_{3},L_{4}=2,4,9,2$ of each stage by default. Stacking the MiM blocks for $L=17$ layers, we build our MiM hierarchical encoder.
We feed the visual sentence embedding outputs of each stage to the respective decoder stages. The spatial shapes of the four-stage encoder outputs are $\frac{H}{8}\times\frac{W}{8}$, $\frac{H}{16}\times\frac{W}{16}$, $\frac{H}{32}\times\frac{W}{32}$ and $\frac{H}{64}\times\frac{W}{64}$, which are different from the feature map scales from typical backbones. To solve this discrepancy, these outputs should pass through straightforward up-sampling modules before reaching the decoder. The module includes a $2\times 2$ transposed convolution with a stride of two, succeeded by batch normalization [46], GeLU [47], a stride-one $3\times 3$ convolution, another batch normalization and GeLU. Therefore our MiM hierarchical encoder can produce feature maps with strides of 4, 8, 16, and 32 pixels relative to the input image.
@@
3.2.4 Complexity Analysis.
Given a visual sequence $T\in\mathbb{R}^{1\times n\times d}$, the computation complexity of SSM is $8ndN$ [34], where $N$ is the SSM dimension and is set to 16 by default. For simplicity, we only consider the core SSM in SS2D and 3 linear layers and calculate the FlOPs of a visual Mamba block as $8ndN+3nd^{2}$. The MiM block includes an inner Mamba block, an outer Mamba block, and a linear layer, with complexities of $8mncN+3mnc^{2}$, $8ndN+3nd^{2}$, and $nmcd$, where $m$ is the number of visual words in a visual sentence and $c$ is the word embedding dimension. Thus, the total FLOPs for the MiM block $FLOPs_{MiM}$ sum to $8mncN+8ndN+3mnc^{2}+3nd^{2}+nmcd$, still maintaining a linear complexity. While the FLOPs of the standard transformer block $FLOPs_{ST}$ is $2nd(6d+n)$ [30]. Considering that $c\ll d$ and $m\ll n$ in high resolution infrared images, the ratio of $FLOPs_{MiM}$ and $FLOPs_{ST}$ approaches 0, meaning that our MiM block introduces a trivial FLOP increase while offering a superior accuracy-efficiency balance demonstrated in subsequent experiments.
@@ 4.1 Experimental Settings
where $A_{i}$ and $A_{u}$ are the areas of intersection and union region between the prediction and ground truth, respectively. $N$ is the total number of samples, $TP[.]$ is the number of true positive pixels, $T[.]$ and $P[.]$ is the number of ground truth and predicted positive pixels.
$P_{d}$ calculates the proportion between the number of correctly predicted targets $N_{pred}$ and all targets $N_{all}$. $F_{a}$ refers to the ratio of falsely predicted target pixels $N_{false}$ and all the pixels in the infrared image $N_{all}$. $P_{d}$ and $F_{a}$ are calculated as follows:
@@ 4.3 Visualization
##\##2 Visual Word $\rightarrow 8\times 8$ 79.50 77.68 100 2.980
##\##3 Visual Sentence $\rightarrow H/4\times W/4$ 79.67 78.04 100 3.685
##\##4 Visual Sentence $\rightarrow H/16\times W/16$ 69.32 67.58 93.52 22.21
##\##5 Ours (VW $\rightarrow 4\times 4$, VS $\rightarrow H/8\times W/8$ ) 80.80 80.20 100 2.179
4.4 Ablation study
4.4.1 Ablation study of Each Module.
4.4.2 Ablation Study of the Granularity of Patch Division.
To explore the impact of information contained in each visual word, we adjust the representation granularity of sub-patches so that each visual word corresponds to a $2\times 2$ (No.#$\#1$) or $8\times 8$ (No.#$\#2$) pixel region in the original image, in contrast to $4\times 4$ of our present setting. We also change the initial spatial shape of visual sentence from $\frac{H}{8}\times\frac{W}{8}$ to $\frac{H}{4}\times\frac{W}{4}$ (No.#$\#3$) and $\frac{H}{16}\times\frac{W}{16}$ (No.#$\#4$) to ablate the effect of visual sentence granularity. We can observe that too large or too small division granularity cannot bring the most ideal performance. We adopt the configuration that demonstrates the best performance in our present setting.